OpenStudy (anonymous):

Write the equation of the line which passes through (5, –2) and is parallel to x = 4.

OpenStudy (anonymous):

use the fomula y-y1=x(x-x1)

OpenStudy (radar):

The line x=4 is a vertical line with a undetermined slope. a parallel line would be another vertical line and passing thru that point would be x=5.

OpenStudy (amistre64):

or see that x=4 is a line that runs straight up and down

OpenStudy (anonymous):

Yea I know that zbay but what do I do with the x=4?

OpenStudy (amistre64):

the equation of a vertical line is just the x part of the point ..

OpenStudy (anonymous):

ohhhhh lol I see, that is why the answer will be in y=mx+b form. Okay thank you guys.

OpenStudy (amistre64):

so the equation to dbl check is; x=5 :)